# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 4 Quantum geometry from adiabatic drives
Having investigated perturbative probes of Virasoro quantum geometry, we now turn to finite deformations $f_{t}$ in the Hamiltonian (2.13). We assume throughout that these deformations are adiabatic and periodic, with period $T$. Then the final state vector $|\psi(T)\rangle$ is nearly proportional to the initial one, $|\psi(0)\rangle$, and one is typically interested in their overlap,
$\langle\psi(0)|\psi(T)\rangle.$ (4.1)
@@
$\gamma_{\textrm{B}}[f_{t}]=\int_{0}^{T}\!\!\oint\frac{\textrm{d}t\,\textrm{d}\varphi}{2\pi}\frac{\dot{f}_{t}(\varphi)}{f_{t}^{\prime}(\varphi)}\Biggl[h-\frac{c}{24}+\frac{c}{24}\biggl(\frac{f_{t}^{\prime\prime}(\varphi)}{f_{t}^{\prime}(\varphi)}\biggr)^{\!\!\prime\,}\Biggr]-\Bigl(h-\frac{c}{24}\Bigr)f_{0}^{-1}(f_{T}(0)),$ (4.2)
where $h$ is the highest weight introduced in (2.16) and $c$ is the central charge. Instead, our goal here is to go beyond leading order and capture the slight misalignment between initial and final states, resulting in a nontrivial return probability given by the squared complex norm of (4.1).
Since the deformations in Eq. (2.11) are implemented by unitary operators in a representation of the group $\operatorname{Diff}(S^{1})$, the return amplitude (4.1) is intimately linked to group theory. This is also true of rotating spins and squeezed states [with groups $\mathrm{SU}(2)$ and $\mathrm{SL}(2,\mathbb{R})$, respectively], indeed of all generalized coherent states whose parameter space is a quotient (2.20). What follows will therefore be phrased, partly, in terms of group manifolds. While we focus for definiteness on $\operatorname{Diff}(S^{1})$, most of our arguments carry over to other groups and coherent states previously studied elsewhere, including e.g. [7, 61, 62].
The plan is as follows. First, we show in general terms how the Schrödinger equation with a time-dependent Hamiltonian (2.11) can be solved in terms of paths in the $\operatorname{Diff}(S^{1})$ group manifold. This expresses the return amplitude (4.1) as an expectation value in a highest-weight state, which ultimately makes it manifest that the return probability is related to the quantum metric. Second, we investigate the specific case of ‘rotating drives’ in $\operatorname{Diff}(S^{1})$. We provide an explicit formula for the resulting displacement between initial and final states, and use this to evaluate the return probability. The consequence, in all cases, is a resonance pattern similar to the one in Fig. 1.
4.1 Return amplitudes from group theory
We begin by showing how the Schrödinger equation with a time-dependent Hamiltonian (2.13) can be solved by time-dependent coherent states of the form (2.17). We then add the assumption of adiabaticity and periodicity to show that the return probability, i.e. the squared norm of (4.1), measures the quantum metric (3.1).
@@ 4.1 Return amplitudes from group theory
Let $f_{t}$ be any family of deformations, and consider the time-dependent Hamiltonian $H(t)\equiv H[f_{t}]$ introduced in Eq. (2.13). We seek to solve the time-dependent Schrödinger equation
$i\partial_{t}|\psi(t)\rangle=H(t)|\psi(t)\rangle=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}|\psi(t)\rangle$ (4.3)
with an initial state $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$, chosen to be an eigenstate of the initial Hamiltonian. [Recall from Eq. (2.16) that $|h\rangle$ is an eigenstate of $H_{0}$ with energy $\tfrac{2\pi}{L}(h-\tfrac{c}{24})$.] The strategy is to reformulate the Schrödinger equation (4.3) as an equation of motion in the group manifold by stripping off information about the Hilbert space and the representation $\cal U$}. To this end, make the ansatz
$|\psi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle,$ (4.4)
where $\alpha_{t}\in\mathbb{R}$ is a time-dependent overall phase (such that $\alpha_{0}=0$) and $g_{t}\in\operatorname{Diff}(S^{1})$ is a path in the group manifold (such that $g_{0}=f_{0}$). Both sides of Eq. (4.3) then take the form of operators acting on the highest-weight state $|h\rangle$. Thus, Eq. (4.3) certainly holds if the operators themselves are equal, i.e. if
$\partial_{t}\alpha_{t}+i\bigl[\partial_{t}\,{\cal U}(g_{t})\bigr]\,{\cal U}(g_{t})^{-1}=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}.$ (4.5)
The right-hand side can be made more explicit thanks to the fact that $H_{0}$ itself is a Lie algebra generator. Indeed, we already computed it in Eq. (2.15). For later convenience, we now write it as ${\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}=\frac{2\pi i}{L}\mathfrak{u}(\mathbf{V}_{t})+C[f_{t}]$, where
@@
is a time-dependent vector field. Note that $V_{t}(\varphi)$ is essentially the dimensionless velocity profile introduced in Eq. (1.1): one has $V_{t}(\varphi)=v_{t}(x)$ for $\varphi=2\pi x/L$.
This is the point where information is needed on the operators ${\cal U}(g)$. As stated in Eq. (2.3), they furnish a projective representation of $\operatorname{Diff}(S^{1})$, so the left-hand side of Eq. (4.5) involves $\partial_{t}\,{\cal U}(g_{t})]\,{\cal U}(g_{t})^{-1}=\mathfrak{u}(\dot{g}_{t}\circ g_{t}^{-1})+ic\partial_{s}\mathsf{C}(g_{s},g_{t}^{-1})\big|_{s=t}$, where $\dot{g}_{t}\equiv\partial_{t}g_{t}$ and $\mathsf{C}(\cdot,\cdot)$ is the Bott cocycle (2.4). We get rid of all terms proportional to the central charge $c$ by choosing the time-dependent phase $\alpha_{t}$ such that
$\partial_{t}\alpha_{t}-c\partial_{s}\mathsf{C}(g_{s},g_{t}^{-1})\big|_{s=t}=C[f_{t}],$ (4.7)
so what remains of Eq. (4.5) is $\mathfrak{u}(\dot{g}_{t}\circ g_{t}^{-1})=\tfrac{2\pi}{L}\mathfrak{u}(\mathbf{V}_{t})$. This, in turn, is satisfied if
$\dot{g}_{t}\circ g_{t}^{-1}=\frac{2\pi}{L}\mathbf{V}_{t},$ (4.8)
which was the desired result: it is an equation for the time-dependent diffeomorphism $g_{t}$ that no longer makes any reference to a Hilbert space or operators thereon. If Eqs. (4.8) and (4.7) hold, then so does the Schrödinger equation (4.3) with the ansatz (4.4).
A straightforward analogue of Eq. (4.8) holds in other situations involving generalized coherent states when the Hamiltonian preserves coherence [i.e. when $H(t)=i\mathfrak{u}(\mathbf{V}_{t})$ for some Lie algebra element $\mathbf{V}_{t}$]. The only difference lies in the choice of group manifold; see e.g. [63, 64] for the case of squeezed states. For $G=\operatorname{Diff}(S^{1})$ as in our current setup, Eq. (4.8) admits a hydrodynamical interpretation: $\mathbf{V}_{t}$ can be seen as a time-dependent vector field that determines the local velocity of a (compressible) fluid on $S^{1}$, and the solution $g_{t}(\varphi)$ of Eq. (4.8) gives the time-dependent position of a fluid parcel initially located at $g_{0}(\varphi)=f_{0}(\varphi)$. Equivalently, $g_{t}$ is the one-parameter flow generated by $\mathbf{V}_{t}$ (with respect to the dimensionless time $2\pi t/L$).
@@
Let us now add two assumptions, namely (i) that the velocity vector field $\mathbf{V}_{t}$ is periodic in time with period $T$, and (ii) that its time dependence is slow, or adiabatic, in the sense that the operator norm of $\partial_{t}H(t)$ is much smaller than $1/L^{2}$. In that limit, the adiabatic theorem states that the solution (4.4) sticks closely to ${\cal U}(f_{t})|h\rangle$, up to a time-dependent phase [65]. Put differently, the coset of $g_{t}$ on the parameter space (2.20) coincides with that of $f_{t}$. The return amplitude (4.1) is thus a pure phase to leading order, involving in particular a Berry phase. As the latter is studied in detail in [53], we do not dwell on it here. Our focus is instead on the first subleading correction to the adiabatic theorem, which accounts for a small difference between the cosets of $g_{t}$ and $f_{t}$ on the parameter space (2.20); see Fig. 4. This leads to a nontrivial return probability given by the square of the amplitude (4.1).
To be specific, let $0<\delta\ll 1$ denote a dimensionless adiabatic parameter, with $\delta\to 0$ indicating the adiabatic limit. The small displacement between the cosets $g_{t}G_{0}$ and $f_{t}G_{0}$ in the parameter space (2.20) can be represented by a time-dependent Lie algebra element $\mathbf{Y}_{t}$ defined as follows: at each time $t$, there exists a finite rotation $R_{\theta(t)}(\varphi)=\varphi+\theta(t)$ by an angle $\theta(t)$ such that
$f_{t}^{-1}\circ g_{t}=R_{\theta(t)}\circ e^{\delta\mathbf{Y}_{t}},$ (4.9)
where $\mathbf{Y}_{t}$ is finite at all times in the adiabatic limit $\delta\to 0$. [Rotations appear here because they span the stabilizer of the undeformed Hamiltonian (2.10).] Using this in the ansatz (4.4), the return amplitude (4.1) can be adiabatically expanded in $\delta\ll 1$ as
$\langle\psi(0)|\psi(T)\rangle$ $=e^{i\alpha_{T}+ic\mathsf{C}(f_{0}^{-1}\!,\,g_{T})}\langle h|\,{\cal U}(f_{0}^{-1}\circ g_{T})|h\rangle$ (4.10)
$=e^{i\Theta}\biggl(1-\frac{\delta^{2}}{2}\Bigl[\langle h|\bigl[i\mathfrak{u}(\mathbf{Y}_{T})\bigr]^{2}|h\rangle-\langle h|i\mathfrak{u}(\mathbf{Y}_{T})|h\rangle^{2}\Bigr]\biggr)+O(\delta^{3}),$ (4.11)
where $\Theta$ is a complicated overall phase—the sum of a dynamical phase and a Berry phase (4.2), plus subleading corrections in $L/T\sim\delta$. For our purposes, the key part of Eq. (4.11) is the one that is not a phase, giving rise to the return probability [7]
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}\sim 1-\delta^{2}\Big(\langle h|\bigl[i\mathfrak{u}(\mathbf{Y}_{T})\bigr]^{2}|h\rangle-\langle h|i\mathfrak{u}(\mathbf{Y}_{T})|h\rangle^{2}\Big)=1-\delta^{2}{\cal G}(\mathbf{Y}_{T},\mathbf{Y}_{T}).$ (4.12)
Here one recognizes the quantum metric (3.1), yielding the squared norm of $\mathbf{Y}_{T}$. [^10] More precisely, the squared norm of $\mathbf{Y}_{T}+\mathfrak{g}_{0}$, seen as a tangent vector of the manifold (2.20) at the identity. What we call $\mathbf{Y}_{T}$ here was called $\mathbf{Y}$ in Eq. (1.6). Note that this is the metric at the identity, even though it measures the distance between the rays of ${\cal U}(g_{T})|h\rangle$ and ${\cal U}(g_{0})|h\rangle$, with $g_{T}$ and $g_{0}$ both well away from the identity. This is again because the parameter space (2.20) is homogeneous, i.e. quantum geometry at the identity determines quantum geometry everywhere.
@@
Figure 4: Cartoon of the parameter space (2.12), showing the periodic drive $f_{t}$ and the corresponding solution $g_{t}$ of Eq. (4.8). In the adiabatic limit, the cosets $f_{t}G_{0}$ and $g_{t}G_{0}$ nearly coincide, but not quite, due to the small micromotion of $g_{t}$. Their mismatch ultimately causes a nontrivial overlap $|\langle\psi(0)|\psi(T)\rangle|=|\langle h|\,{\cal U}(g_{0})^{-1}\,{\cal U}(g_{T})|h\rangle|\neq 1$ that essentially measures the quantum distance between nearby Gaussian wave functions. See also Fig. 5 below.
4.2 Return probabilities for rotating drives

### v2

@@ IV Adiabatic quantum geometry
Having investigated perturbative probes of Virasoro quantum geometry, we now turn to finite deformations $f_{t}\in\operatorname{Diff}(S^{1})$ in the Hamiltonian (19). We assume throughout that these deformations are adiabatic and periodic, with period $T$ so that $H[f_{T}]=H[f_{0}]$. [^12] Beware that this does not imply $f_{T}=f_{0}$, but only $f_{T}(\varphi)=f_{0}(\varphi+\theta)$ for some angle $\theta$. Then the final state vector $|\psi(T)\rangle$ is nearly proportional to the initial one, $|\psi(0)\rangle$, and one is typically interested in their overlap,
$\langle\psi(0)|\psi(T)\rangle.$ (44)
@@
$\gamma_{\textrm{B}}[f_{t}]$ $=\int_{0}^{T}\!\!\!\textrm{d}t\oint\frac{\textrm{d}\varphi}{2\pi}\frac{\partial_{t}f_{t}(\varphi)}{f_{t}^{\prime}(\varphi)}\Biggl[h-\frac{c}{24}+\frac{c}{24}\biggl(\!\frac{f_{t}^{\prime\prime}(\varphi)}{f_{t}^{\prime}(\varphi)}\!\biggr)^{\!\!\prime\,}\Biggr]$
$\quad-\Bigl(h-\frac{c}{24}\Bigr)\,f_{0}^{-1}(f_{T}(0)),$ (45)
where $h$ is the highest weight introduced in (22) and $c$ is the central charge. Instead, our goal here is to go beyond leading order and capture the slight misalignment between initial and final states, resulting in a nontrivial return probability given by the squared complex norm of (44).
Since the deformations in Eq. (17) are implemented by unitary operators in a representation of the group $\operatorname{Diff}(S^{1})$, the return amplitude (44) is intimately linked to group theory. This is also true of rotating spins and squeezed states [with groups $\mathrm{SU}(2)$ and $\mathrm{SL}(2,\mathbb{R})$, respectively], indeed of all generalized coherent states whose parameter space is a quotient (26) [84, 67]. What follows will thus be phrased, partly, in terms of group manifolds. While we focus for definiteness on $\operatorname{Diff}(S^{1})$, most of our arguments carry over to other groups and coherent states previously studied elsewhere, including e.g. [6, 92, 93].
The plan is as follows. First, we show in general terms how the Schrödinger equation with a time-dependent Hamiltonian (17) can be solved in terms of paths in the $\operatorname{Diff}(S^{1})$ group manifold. This expresses the return amplitude (44) as an expectation value in a highest-weight state, which ultimately makes it manifest that the return probability is related to the quantum metric (28). Second, we investigate the specific case of ‘rotating drives’ in $\operatorname{Diff}(S^{1})$. We provide an explicit formula for the resulting displacement between initial and final states, and use this to evaluate the return probability. The consequence, in all cases, is a resonance pattern like the one in Fig. 1.
IV.1 Return amplitudes from group theory
We begin by showing how the Schrödinger equation with a time-dependent Hamiltonian (19) can be solved by time-dependent coherent states of the form (23). We then add the assumption of adiabaticity and periodicity to show that the return probability, i.e. the squared complex norm of (44), probes the quantum metric (28).
@@ IV.1 Return amplitudes from group theory
Let $f_{t}$ be any family of deformations, and consider the time-dependent Hamiltonian $H(t)\equiv H[f_{t}]$ introduced in Eq. (19). We seek to solve the time-dependent Schrödinger equation
$i\partial_{t}|\psi(t)\rangle=H(t)|\psi(t)\rangle=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}|\psi(t)\rangle$ (46)
with an initial state $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$, chosen to be an eigenstate of the initial Hamiltonian. [Recall from Eq. (22) that $|h\rangle$ is an eigenstate of $H_{0}$ with energy $\tfrac{2\pi}{L}(h-\tfrac{c}{24})$.] The strategy is to reformulate the Schrödinger equation (46) as an equation of motion in the group manifold by stripping off information about the Hilbert space and the representation $\cal U$}. To this end, make the ansatz
$|\psi(t)\rangle=e^{i\alpha_{t}}\,{\cal U}(g_{t})|h\rangle,$ (47)
where $\alpha_{t}\in\mathbb{R}$ is a time-dependent overall phase (such that $\alpha_{0}=0$) and $g_{t}\in\operatorname{Diff}(S^{1})$ is a path in the group manifold (such that $g_{0}=f_{0}$). Both sides of Eq. (46) then take the form of operators acting on the highest-weight state $|h\rangle$. Thus, Eq. (46) certainly holds if the operators themselves are equal, i.e. if
$\partial_{t}\alpha_{t}+i\bigl[\partial_{t}\,{\cal U}(g_{t})\bigr]\,{\cal U}(g_{t})^{-1}=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}.$ (48)
The right-hand side can be made more explicit thanks to the fact that $H_{0}$ itself is a Lie algebra generator. Indeed, we already computed it in Eq. (21). For later convenience, we now write it as ${\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}=\frac{2\pi i}{L}\mathfrak{u}(\mathbf{V}_{t})+C[f_{t}]$, where
@@
is a time-dependent vector field. Note that $V_{t}(\varphi)$ is essentially the dimensionless velocity profile introduced in Eq. (1): one has $V_{t}(\varphi)=v_{t}(x)$ for $\varphi=2\pi x/L$.
This is where information on the operators ${\cal U}(g)$ is needed. As stated in Eq. (9), they furnish a projective representation of $\operatorname{Diff}(S^{1})$, so the left-hand side of Eq. (48) involves $\partial_{t}\,{\cal U}(g_{t})]\,{\cal U}(g_{t})^{-1}=\mathfrak{u}(\partial_{t}g_{t}\circ g_{t}^{-1})+ic\partial_{s}\mathsf{C}(g_{s},g_{t}^{-1})\big|_{s=t}$, where $\mathsf{C}(\cdot,\cdot)$ is the Bott cocycle (10). We get rid of all terms proportional to the central charge $c$ by choosing the time-dependent phase $\alpha_{t}$ such that
$\partial_{t}\alpha_{t}-c\partial_{s}\mathsf{C}(g_{s},g_{t}^{-1})\big|_{s=t}=C[f_{t}],$ (50)
so what remains of Eq. (48) is $\mathfrak{u}(\partial_{t}g_{t}\circ g_{t}^{-1})=\tfrac{2\pi}{L}\mathfrak{u}(\mathbf{V}_{t})$. This, in turn, is satisfied if
$\partial_{t}g_{t}\circ g_{t}^{-1}=\frac{2\pi}{L}V_{t}.$ (51)
The latter is an equation for the time-dependent diffeomorphism $g_{t}$ that no longer makes any reference to a Hilbert space or operators thereon. If Eqs. (51) and (50) hold, then so does the Schrödinger equation (46) with the ansatz (47).
A straightforward analogue of Eq. (51) holds in other situations involving generalized coherent states when the Hamiltonian preserves coherence [i.e. when $H(t)=i\mathfrak{u}(\mathbf{V}_{t})$ for some Lie algebra element $\mathbf{V}_{t}$]. The only difference lies in the choice of group manifold; see e.g. [94, 95] for the case of squeezed states. For $G=\operatorname{Diff}(S^{1})$ as in our current setup, Eq. (51) admits a hydrodynamical interpretation: $\mathbf{V}_{t}$ can be seen as a time-dependent vector field that determines the local velocity of a (compressible) fluid on $S^{1}$, and the solution $g_{t}(\varphi)$ of Eq. (51) gives the time-dependent position of a fluid parcel initially located at $g_{0}(\varphi)=f_{0}(\varphi)$. Equivalently, $g_{t}$ is the one-parameter flow generated by $\mathbf{V}_{t}$ (with respect to the dimensionless time $2\pi t/L$).
@@
Let us now add two assumptions, namely (i) that the velocity vector field $\mathbf{V}_{t}$ is periodic in time with period $T$, and (ii) that its time dependence is slow, or adiabatic, in the sense that the operator norm of $\partial_{t}H(t)$ is much smaller than $1/L^{2}$. In that limit, the adiabatic theorem states that the solution (47) sticks closely to ${\cal U}(f_{t})|h\rangle$, up to a time-dependent phase [96]. Put differently, the coset of $g_{t}$ on the parameter space (26) coincides with that of $f_{t}$. The return amplitude (44) is thus a pure phase to leading order, involving in particular a Berry phase. As the latter is studied in detail in [46], we do not dwell on it here. Our focus is instead on the first subleading correction to the adiabatic theorem, which accounts for a small difference between the cosets of $g_{t}$ and $f_{t}$ on the parameter space (26); see Fig. 4. This leads to a nontrivial return probability given by the square of the amplitude (44).
@@
Figure 4: Cartoon of the parameter space (26), with $G=\operatorname{Diff}(S^{1})$ and $G_{0}=S^{1}$ as in Eq. (18), showing the periodic drive $f_{t}$ and the corresponding solution $g_{t}$ of Eq. (51). In the adiabatic limit, the cosets $f_{t}G_{0}$ and $g_{t}G_{0}$ nearly coincide, but not quite, due to the small micromotion of $g_{t}$. Their mismatch ultimately causes a nontrivial overlap $|\langle\psi(0)|\psi(T)\rangle|=|\langle h|\,{\cal U}(g_{0})^{-1}\,{\cal U}(g_{T})|h\rangle|\neq 1$ that measures the quantum distance between nearby coherent states. See also Fig. 5 below.
To be specific, let $0<\delta\ll 1$ denote a dimensionless adiabatic parameter, with $\delta\to 0$ indicating the adiabatic limit. [The exact value of $\delta$ depends on the path $f_{t}$, but it is safe in general terms to think of it as $\delta\sim L/T$; see e.g. Eq. (60).] Let $G_{0}\cong S^{1}$ be the stabilizer that leaves the undeformed CFT Hamiltonian invariant, as in Eq. (18). The small displacement between the cosets $g_{t}G_{0}$ and $f_{t}G_{0}$ in the parameter space (26) can then be represented by a time-dependent Lie algebra element $\mathbf{Y}_{t}$ defined as follows: at each time $t$, there exists a finite rotation $R_{\theta(t)}(\varphi)=\varphi+\theta(t)$ by an angle $\theta(t)$ such that
$f_{t}^{-1}\circ g_{t}=R_{\theta(t)}\circ e^{\delta\mathbf{Y}_{t}},$ (52)
where $\mathbf{Y}_{t}$ is finite at all times in the adiabatic limit $\delta\to 0$. [Rotations appear here because they span the stabilizer of the undeformed Hamiltonian (16).] Using this in the ansatz (47), the return amplitude (44) can be adiabatically expanded in $\delta\ll 1$ as
$\langle\psi(0)|\psi(T)\rangle=e^{i\alpha_{T}+ic\mathsf{C}(f_{0}^{-1}\!,\,g_{T})}\langle h|\,{\cal U}(f_{0}^{-1}\circ g_{T})|h\rangle$
$\sim e^{i\Theta}\biggl[1-\frac{\delta^{2}}{2}\Bigl(\langle h|\bigl[i\mathfrak{u}(\mathbf{Y}_{T})\bigr]^{2}|h\rangle-\langle h|i\mathfrak{u}(\mathbf{Y}_{T})|h\rangle^{2}\Bigr)\biggr]$
$=e^{i\Theta}\biggl[1-\frac{\delta^{2}}{2}{\cal G}(\mathbf{Y}_{T},\mathbf{Y}_{T})\biggr],$ (53)
up to $O(\delta^{3})$ corrections. Here, $\Theta$ is an overall phase—the sum of a dynamical phase, a Berry phase (45), and subleading corrections in $L/T\sim\delta$. For our purposes, the key part of Eq. (53) is the one that is not a phase, involving the squared quantum norm (28) of $\mathbf{Y}_{T}$. [^14] More precisely, the squared norm of $\mathbf{Y}_{T}+\mathfrak{g}_{0}$, seen as a tangent vector of the manifold (26) at the identity. This gives rise to the return probability [6]
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}=1-\delta^{2}{\cal G}(\mathbf{Y}_{T},\mathbf{Y}_{T})+O(\delta^{3})$ (54)
expressed in terms of the quantum metric, as announced in Eq. (6). Note that this is the metric at the identity, even though it measures the distance between the rays of ${\cal U}(g_{T})|h\rangle$ and ${\cal U}(g_{0})|h\rangle$, with $g_{T}$ and $g_{0}$ both well away from the identity. This is again because the parameter space (26) is homogeneous: its quantum geometry at the identity determines quantum geometry everywhere.
IV.2 Return probabilities for rotating drives